OpenStudy (anonymous):

need help finding intervals of increse....max and min...intervals of concavity up and down...and point of inflection f '(x)=3e^x/(3+e^x)^2

OpenStudy (amistre64):

so this is our y' right?

OpenStudy (amistre64):

e^x never equals zero, so there are no critical points

OpenStudy (anonymous):

ok but i get an interval of increase right from some number to infinity

OpenStudy (amistre64):

I would say yes, but to check for concavity we tend to use the 2nd derivative right?

OpenStudy (anonymous):

well i know for sure that my function is increasing so there has to be interval of increasing but i dont know the starting point

OpenStudy (amistre64):

e^x is never a negative number, so all of its positive right?

OpenStudy (anonymous):

figured it out it from -infinity to infinity for increase variables ok now for concavity...

OpenStudy (amistre64):

concave up is my guess.... but use 2nd derivative to be certain

OpenStudy (anonymous):

i meant increase interval is -infinity to infinity

OpenStudy (amistre64):

(3e^x) (3+e^x)^-2 -6 e^2x (3e^x) --------- + ----------- right? (3+e^x)^3 (3+e^x)^2

OpenStudy (amistre64):

3 e^2x --------- (3+e^x)^3

OpenStudy (amistre64):

its always concave up....

OpenStudy (anonymous):

my sources say otherwise and i am looking for the intervals from which it is concave up and down

OpenStudy (amistre64):

i mighta goofed on the 2nd integral.... better double check

OpenStudy (anonymous):

\[f"(x)=-(3e^{2x}-9e^x)/e ^{3x}+9e ^{2x}+27e^x+26\]

OpenStudy (amistre64):

3e^x( e^3x -2e^2x +9e^x ) ------------------------ (3+e^x)^4

OpenStudy (anonymous):

no it not to the fourth power...

OpenStudy (amistre64):

havent expanded the bottom to see if anything cancels, but thats just a left over from the quotient rule

OpenStudy (amistre64):

it either that, or my heads got all jumbled..... which is very likely :)

OpenStudy (anonymous):

oh no i think things cancel i am just saying the function i have is the correct function already checked it ....lol

OpenStudy (amistre64):

my top gets to be: -3e^x (e^x -1) i cant get it to match yours

OpenStudy (amistre64):

-3e^x (e^x -3)

OpenStudy (amistre64):

now it does lol

OpenStudy (anonymous):

yea there you go so x=ln 3

OpenStudy (amistre64):

yeah, when e^x = 3 x = ln(3) thats right... we get an inflection right?

OpenStudy (amistre64):

everything to the left of that would be (+) and everything to the right of it would be(-)

OpenStudy (anonymous):

yea...problem one complete now for two....consider the function f(x)=2x+4/3x+2 for this function there are two important intervals (-infinity, A) and (A, infinity) where the function is not defined at A. Find A

OpenStudy (amistre64):

leftside is concave up, and rightside of it os concave down

OpenStudy (amistre64):

A = -2/3 there is a vertical asymptote there and possess no real value.

OpenStudy (amistre64):

nothing calcels top to bottom, so it aint a hole; its just the place where the denominator goes to zero... -2/3

OpenStudy (anonymous):

ok yea that was correct....an i see you clarification

OpenStudy (anonymous):

problem 3 sin^2(x/5) defined at interval [-14.907963,3.226991]

OpenStudy (anonymous):

find where f(x) is concave down, global mini of the function, a local max which is not a global max, the function is increasing on a region....note: some answers must be given in interval notation

OpenStudy (amistre64):

the normal period for pi is about 6 and this stretches it past 6 and goes to 10pi for the period. 10pi is about 31.4

OpenStudy (amistre64):

the amplitude is exaggerated, but thats just the y value....

OpenStudy (amistre64):

sin normally has a max at pi/2 and a min at 3pi/2 so, this one will be at 5pi/2 and 15pi/2 if I am picturing it correctly

OpenStudy (amistre64):

but lets work it :)

OpenStudy (amistre64):

Dx (3 sin^2(x/5)) y' = (3/5) sin(x/5)cos(x/5)

OpenStudy (amistre64):

if sin or cos are zero, we get zeros...

OpenStudy (anonymous):

the fx is sin^2(x/5) sorry its problem 3....

OpenStudy (amistre64):

sin = 0 at 0 and 180, cos = 0 at 90 and 270

OpenStudy (amistre64):

replace (3/5) with (1/5) same setup tho

OpenStudy (amistre64):

but because the normal period is elongated to 10pi, we gotta find those numbers :)

OpenStudy (amistre64):

they are..... 0, 5pi, 5pi/2, 15pi/2

OpenStudy (amistre64):

we can go backwards to... -5pi, -5pi/2, -15pi/2, -10pi

OpenStudy (amistre64):

which of these numbers is in our interval?

OpenStudy (amistre64):

5pi is bigger than 15

OpenStudy (amistre64):

[-14.907963 , 3.226991] -15/2 = -7.5, so -5pi/2 is in our interval 0 is in our interval we got these 2 numbers (-5pi/2) and (0)

OpenStudy (anonymous):

are you using \[\sin ^2(x/5)\]

OpenStudy (amistre64):

yes, I am using the derivative of that and checking for zeros :)

OpenStudy (amistre64):

the 3 was superfluous and could be tossed on a whim...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok the global mini is at 0 but now we need a local max

OpenStudy (amistre64):

-5pi/2 would be the local max...

OpenStudy (anonymous):

nah it not...i checked

OpenStudy (amistre64):

this is what we look like I think

OpenStudy (anonymous):

our local max was 3.22 from the interval

OpenStudy (anonymous):

i need the interval where it is concave down and region where it is increasing

OpenStudy (amistre64):

computer froze